OpenStudy (anonymous):

a rectangular garden has a perimeter of 48 cm and an area of 140 sq. cm. what is the length of this garden

OpenStudy (mathmath333):



OpenStudy (mathmath333):

\(\Large a+b+a+b=48\\ \Large 2a+2b=48\\ \Large a+b=24---1\)

OpenStudy (mathmath333):

\(\Large area~ of~ ractangle=140\\ \Large length \times ~breadth=140\\ \Large a\times b=140--------2\)